OpenStudy (anonymous):

Find the cooordinates of the x and y intercepts of: 2x + 3y = -21

OpenStudy (anonymous):

it asks for numerator and denominator of x intercept an then ________ for the other

OpenStudy (anonymous):

-23/3?

OpenStudy (anonymous):

reduce to the form x/a +y/b =1

OpenStudy (anonymous):

2x + 3y = -21 (2x/-21) +(3y/-21) =1

OpenStudy (anonymous):

(x/(-21/2) ) + ( y/ (-21/3) ) =1

OpenStudy (anonymous):

so xintercept is (-21/2) and y-intercept is (-21/3)

OpenStudy (anonymous):

it asks for a fraction?

OpenStudy (anonymous):

cordinates will be ( (-21/2) ,0 ) and (0, (-21/3) )

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

it wont let me put a fraction in the coordinate

OpenStudy (anonymous):

see hold on

OpenStudy (anonymous):

Is this a test your taking?

OpenStudy (anonymous):

no its a practice quiz online i told u

OpenStudy (anonymous):

before